Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Puzzling probability question

Expert replies
by spucmu » Sat Apr 30, 2011 12:40 pm
Attached is the probability question.

Shouldn't answer to this question be B?

There are two possibilities as I can see:
30/500 * 1/800 + 30/800 * 1/500

Thanks,
Swapnil
You do not have the required permissions to view the files attached to this post.
Join the discussion
Source: — Problem Solving |

by edvhou812 » Sat Apr 30, 2011 1:19 pm
This is an AND probability problem. Formula A*B
So you need to multiply two numbers to get your answer.

First Number (Probability of selecting a business student who has a sibling in the law school):
There are 500 business students, of which 30 have siblings in the law school. Probability of randomly selecting a business student who has a sibling at the law school is 30/500

There are 800 students in the law school, but only one can be a sibling of the business student we randomly selected. Probability of random selection is 1/800

Probability that we randomly select a sibling pair from the business school and law school:

30/500*1/800 = 3/40,000
Join the discussion

by spucmu » Sat Apr 30, 2011 3:53 pm
Shouldn't this should be AND & OR problem.

(Probability of selecting a business student who has a sibling in the law school) OR (Probability of selecting a law student who has a sibling in the business school) =
30/500 * 1/800 + 30/800 * 1/500
Join the discussion

by edvhou812 » Sat Apr 30, 2011 3:59 pm
spucmu wrote:Shouldn't this should be AND & OR problem.

(Probability of selecting a business student who has a sibling in the law school) OR (Probability of selecting a law student who has a sibling in the business school) =
30/500 * 1/800 + 30/800 * 1/500
Not an And/Or problem. Just And. If you wanted to select the law student first and then the business school student, the formula would be 30/800*1/500 which still yields the same answer 3/40000
Join the discussion

by GMATGuruNY » Sun May 01, 2011 1:54 am
A certain business school has 500 students, and the law school at the same university has 800 students. Among these students, there are 30 sibling pairs consisting of 1 business student and 1 law student. If 1 student is selected at random from both schools, what is the probability that a sibling pair is selected?
It might be easier to look at the problem this way:

A pair is to be formed by choosing 1 student from the business school and 1 student from the law school.
Number of choices from the business school = 500.
Number of choices from the law school = 800.
Total number of ways to choose a pair = 500*800 = 400,000.

Of these 400,000 pairs, the number of sibling pairs = 30.

P(sibling pair) = 30/400,000 = 3/40,000.
Private tutor exclusively for the GMAT and GRE, with over 20 years of experience.
Followed here and elsewhere by over 1900 test-takers.
I have worked with students based in the US, Australia, Taiwan, China, Tajikistan, Kuwait, Saudi Arabia -- a long list of countries.
My students have been admitted to HBS, CBS, Tuck, Yale, Stern, Fuqua -- a long list of top programs.

As a tutor, I don't simply teach you how I would approach problems.
I unlock the best way for YOU to solve problems.

For more information, please email me (Mitch Hunt) at [email protected].
Student Review #1
Student Review #2
Student Review #3
Join the discussion

by dineshkundu » Tue May 03, 2011 10:45 pm
Hi GMATGuruNY,

I also faced this problem while giving a test, and i did the same mistake as mentioned in post above, I treated the question in similar fashion : -
(Probability of selecting a business student who has a sibling in the law school) OR (Probability of selecting a law student who has a sibling in the business school)
But obviously i was wrong.
How to decide when using to use OR and when to use AND. I certainly could not figure this out on this question.

Thanks
Dinesh
Join the discussion

by Brent@GMATPrepNow » Thu Jun 06, 2019 6:37 am
spucmu wrote:Attached is the probability question.
P(selecting a sibling pair) = P(select a business student with a sibling AND select a law student who is that business student's sibling)
= P(select a business student with a sibling) x P(select a law student who is that business student's sibling)
= 30/500 x 1/800
= 30/400,000
= 3/40,000
= A

Note: P(select a business student with a sibling) = 30/500, because 30 of the 500 business students have a sibling in law school.
P(select a law student who is that business student's sibling) = 1/800, because there are 800 law students and only 1 is the sibling of the selected business student.

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

by Scott@TargetTestPrep » Mon Jun 10, 2019 6:06 pm
spucmu wrote:Attached is the probability question.
The probability of selecting from the business school any one sibling from the 30 sibling pairs is 30/500. Once that person is selected, the probability of selecting his or her sibling from the law school is 1/800; thus, the probability of selecting a sibling pair is:

30/500 x 1/800 = 3/50 x 1/800 = 3/40000

Alternatively, the probability of selecting from the law school any one sibling from the 30 sibling pairs is 30/800. Once that person is selected, the probability of selecting his or her sibling from the business school is 1/500; thus, the probability of selecting a sibling pair is:

30/800 x 1/500 = 3/80 x 1/500 = 3/40000

Answer: A

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion